### v2

# DyFFPAD: Dynamic Fusion of Convolutional and Handcrafted Features for Fingerprint Presentation Attack Detection

Abstract
Automatic fingerprint recognition systems suffer from the threat of presentation attacks due to their wide range of deployment in areas including national borders and commercial applications. A presentation attack can be performed by creating a spoof of a user’s fingerprint with or without their consent. This paper presents a dynamic ensemble of deep CNN and handcrafted features to detect presentation attacks in known-material and unknown-material protocols of the livness detection competition. The proposed presentation attack detection model, in this way, utilizes the capabilities of both deep CNN and handcrafted features techniques and exhibits better performance than their individual performances. The proposed method is validated using benchmark databases from the Liveness Detection Competition in 2015, 2017, and 2019, yielding overall accuracy of 96.10%, 96.49%, and 94.99% on them, respectively. The proposed method outperforms state-of-the-art methods in terms of classification accuracy.
Index Terms:
Presentation Attack Detection, Hybrid Architecture, Handcrafted Features, Deep CNN, Fingerprint Biometrics.
1 Introduction
Automatic Fingerprint Recognition Systems (AFRS) are most widely used for the purpose of person authentication and verification. Their user-friendliness and cost-effectiveness make them popular while validating the identity of persons at airports, national borders, and distribution of government-funded aid. The utilization of these systems in a wide range of applications, also makes them vulnerable to some internal and external security threats. Presentation Attack (PA), or spoofing, is an external attack where an artificial artifact of a user’s finger is presented to the sensor of an Automated Fingerprint Recognition System (AFRS). PAs or spoofs can be created either by a cooperative method or a non-cooperative method of spoofing. Fabrication materials such as latex, woodglue, gelatine, silicon, ecoflex, etc. are available at a very low cost to create the spoof of a fingerprint. Fingerprint Presentation Attack Detection (FPAD) serves as a countermeasure to these attacks and to empower an AFRS against them. The recent FPAD methods suggested by various researchers are categorized into two broad classes including hardware-based and software-based methods. Hardware-based methods are quite expensive due to the utilization of additional sensing devices that measure natural properties such as heart rate, odor, temperature, etc. The utilization of these sensors makes hardware-based methods less user-friendly and quite expensive for an organization to use in access-control applications. On the other side, software-based methods require only fingerprint samples which makes them friendly to the user and cost-effective to the organization as compared to hardware-based methods. Current software-based methods can be further categorized into perspiration and pore based-methods [9], [2] statistical and handcrafted feature-based methods [32], [31], [37], [3] and deep-learning based methods [5], [4], [28]. Perspiration-based methods suffer from dryness in the environment which sometimes causes the rejection of a genuine fingerprint sample. Similarly, pore-based methods necessitate the sensing device to capture a high-definition image of the fingertip. This necessity impacts the overall cost of these methods. Existing works utilizing statistical and handcrafted feature extract some predefined features from the input fingerprint sample to classify them as live or spoof. These methods are affected by the quality of the fingerprint samples. However, some of these methods including Sharma et al.[32], Sharma et al. [31] have shown good FPAD capability while the spoof sample is created using known material but they have not been tested properly on the spoofs fabricated with unknown materials. Deep learning methods consist of a classifier that possesses convolutional layers as a feature extractor. These layers have an unmatched capability of extracting minute features from the input samples. In recent times, various methods based on deep learning have been suggested by various researchers which have exhibited better performance than other software-based methods but they also suffer while being tested with the spoofs fabricated with unknown materials.
In our proposed work, a dynamic ensemble of handcrafted and deep features based model is presented. This Dynamic Fusion of convolutional and handcrafted features for Fingerprint Presentation Attack Detection (DyFFPAD), is an end-to-end approach that dynamically incorporates deep and handcrafted features. The model is build using two sub-models that work in collaboration to classify the live and spoof fingerprint samples. The first sub-model is a feature extractor model comprising of Deep Neural Network (DNN) which is empowered with an image descriptor namely Local Phase Quantization (LPQ) and a set of quality-based handcrafted features including Ridge Valley Clarity (RVC), Gabor quality, Frequency Domain Analysis (FDA), and Orientation Flow (OFL), etc. On the other hand, the second sub-model is a DenseNet architecture’s feature base attached to a custom-made DNN. DenseNet has shown remarkable performance while being tested on some well-known image classification problem databases including MNIST [7], CIFAR [36], and imagenet [6]. The output of both the sub-models is combined and fed to another DNN which has a single neuron in the last for the binary classification. The TSNE plots depicted in Fig. 1 show that the proposed method is able to distinguish between live samples, and spoof samples irrespective of their fabrication materials. The proposed model exhibits the same performance on different datasets and its efficacy does not affect whether the spoofs are fabricated with known or unknown materials. The performance of the proposed method is validated using Liveness Detection Competition (LivDet) databases, demonstrating superior results compared to state-of-the-art methods in both intra-sensor same-material and intra-sensor cross-material protocols. A detailed comparison with the state-of-the-art methods indicates the supremacy of the proposed method while detecting the PAs in benchmark protocols of FPAD.
The main contributions of this paper are highlighted below.
1. A novel end-to-end architecture is proposed that embodies the capabilities of convolutional filters along with LPQ and handcrafted features to detect PAs fabricated using known and unknown materials.
2. A dynamic fusion of handcrafted and convolutional features is proposed which addresses the limitations that occur in static fusion.
3. An experimental study is presented that shows the performance of PAs on the DenseNet classifier and DNNs which are trained with handcrafted and LPQ features individually and a combined feature vector comprised of LPQ and handcrafted feature values.
4. An exhaustive comparative study, conducted on intra-sensor known-material and unknown-material protocols, reveals that the proposed method outperforms others in terms of classification accuracy.
5. From TSNE representations, it is evident that the proposed model demonstrates consistency while detecting the spoofs fabricated with known and unknown materials.
The remainder of this paper is organized as follows: Section 2 reviews state-of-the-art methods proposed by researchers for the detection of PAs, highlighting their advantages and limitations. Section 3 details the design and functionality of the proposed architecture. Section 5 presents the experimental results, followed by a comparative analysis. Finally, the conclusion is provided in Section 6.
(a) LivDet 2015 Biometrika                                               (c) LivDet 2015 Crossmatch
(b) LivDet 2015 Digital Persona                                             (d) LivDet 2015 Greenbit
Figure 1: 2D - tsne plots for the features generated by the proposed DyFFPAD model on LivDet 2015 (a) Biometrika, (b) Digital Persona, (c) Crossmatch, and (d) Greenbit sensors. 0 represents spoofs fabricated with known materials, 1 represents spoofs fabricated with unknown materials, and 2 represents live fingerprint samples. It depicts the proposed model’s ability for generalization across both known and unknown materials in fabricated spoofs, effectively distinguishing them from live fingerprint samples.
2 Related Work
PAs represent the most significant security threat to the AFRS The applications of AFRS in security and commercial areas brought the attention of researchers toward the development of cost-effective and user-friendly solutions to this problem. In this section, methods put forth by researchers to keep the AFRS safe against PAs are discussed. Existing methods utilizes perspiration and pore features, statistical and handcrafted features, and deep learning features. The methods that fall into these categories and their limitations are discussed in this section.
2.1 Perspiration and pore based-methods
Perspiration is a natural property in human skin that is caused by the presence of tiny holes or pores. As spoofs created with different materials do not have this natural property, it could be utilized to develop a method that distinguishes between live and spoof fingerprints based on this property. In this attempt, Derakshini et al. [8] leveraged the presence of perspiration and sweat diffusion patterns to differentiate between live and spoofed fingerprints. Further, Abhyankar et al. [2] proposed a wavelet-based approach to detect PAs using the fingerprint’s perspiration property. A custom-made fingerprint database is utilized to validate the performance of the proposed method. The count of pores within a live fingerprint and its counterfeits crafted from various materials can exhibit variability because pores are hard to reflect in the spoofs during the fabrication process. Espinoza [9] utilized this dissimilarity as a distinguishing feature to detect PAs. The proposed method is validated on a custom-made dataset. Marcialis et al. [21] also proposed a similar method that detects pores from multiple fingerprints captured at an interval of five seconds. The number of pores is further in the subsequent samples is further utilized as a feature to detect PAs. The proposed method is validated on a custom-made database having 8,960 live and 7,760 spoof samples. Manivanan et al. [20] utilized high-pass filtering followed by correlation filtering techniques [22] to identify active sweat pores on fingerprint ridges. The results demonstrate distinguishable correlation peaks corresponding to the sweat pores on the ridges. Regardless of how the perspiration pattern is employed to detect PAs, its presence is dependent on the temperature of the surrounding environment. Sometimes, even a live finger does not exhibit this property in a dry environment which occasionally prompts the FPAD method to mis-classify this live sample as a spoof. Similarly, the pore-based methods cost high due to the utilization of a high-resolution sensing device that is capable of collecting fingerprint samples with higher dpi (>=>=> =1000 pixels per inch). These factors collectively make perspiration-based and pore-based methods less economical and less convenient to use.
2.2 Statistical and handcrafted feature based-methods
The inherent disparities between the natural attributes of a finger’s skin and counterfeit spoofs, such as variations in color and moisture levels, manifest in the quality of collected samples. This phenomenon prompted researchers to utilize the quality of captured fingerprint samples as a distinguishing feature. This section describes some of the approaches that fall within this category. Park et al. [27] employed a range of statistical features, encompassing deviation, hyper-flatness, average brightness, standard deviation, variance, skewness, hyper-skewness, kurtosis and differential image features, for the training of a Support Vector Machine (SVM) classifier, enabling detection of live and spoof samples. The efficacy of the proposed approach is evaluated through experimentation on the ATVSFFp database, comprising 272 genuine and 270 spoof fingerprint samples. In a similar attempt, Xia et al. [37] proposed an image descriptor that captures both intensity variance and gradient properties to create a combined feature vector. This feature vector is then used to train an SVM classifier for spoof detection. The proposed method is validated using the LivDet 2011, 2013, and 2015 databases. Kim et al. [18] proposed an image descriptor that utilizes the local coherence of fingerprints as a feature to train an SVM classifier. LivDet 2009, 2011, 2013, and 2015 and ATVSFFp databases are utilized by the authors to validate their work. Yuan et al. [40] introduced a novel approach that utilizes gradient properties for image analysis. This method involves generating image gradient matrices through diverse quantization operators which are derived from the computation of two co-occurrence matrices using the Laplacian operator. These computed matrices are then employed to train a back-propagation neural network. The effectiveness of the proposed method is demonstrated using the LivDet 2013 database. Gragnaniello et al. [13] employed the Weber Local Descriptor (WLD) to extract digital excitation and gradient information from fingerprint samples. These extracted features are subsequently used to train an SVM classifier. The proposed method is validated using the LivDet 2009 and 2011 databases. In another work, Gragnaniello et al. [14] proposed a novel image descriptor named Local Contrast Phase Descriptor (LCPD) which is based on a component of spatial domain and rotation invariant version of Local Phase Quantization (LPQ). The proposed method is validated on LivDet 2011 database. Rattani et al. [29] proposed an incremental learning based approach that consist of a novel material detector along with the classifier. This detector classifies input fingerprint samples into three categories: live, spoof, and unknown. Samples classified as ’novel’ are further used to train the classifier. The proposed method is validated using LivDet 2011 database. In continuation to their previous work [29], Rattani et al. [30] proposed a novel approach that employs a Weibull-calibrated SVM for the classification of fingerprint samples. This approach integrates one-class and binary SVMs, demonstrating a significant improvement over their previous method [29]. Since the finger skin and spoof fabrication materials exhibit different levels of elasticity the resultant fingerprints have some differences in terms of quality. The difference in the live and spoof samples fabricated with different materials results in uneven ridge and valley widths and the fabrication process introduces other quality irregularities in the spoofs. In an attempt to develop a quality-based FPAD model, Sharma et al. [32] utilized some fingerprint quality features for the detection of PAs. They extracted Ridge and Valley Smoothness (RVS), Ridge and Valley Clarity (RVC), Frequency Domain Analysis (FDA), and Orientation Certainty Level (OCL) which are combined to form a feature vector. The feature vector is subsequently utilized to train a Random-Forest (RF) classifier. The proposed method is validated on a variety of databases including LivDet 2009, 2011, 2013, and 2015. Similarly, Sharma et al. [31] proposed a novel variant of the Local Binary Pattern (LBP) known as the Local Adaptive Binary Pattern (LABP). This feature is combined with the existing Complete Local Binary Pattern (CLBP) and Binary Statistical Image Features (BSIF) to train both an SVM classifier and a Deep Neural Network (DNN). The effectiveness of the proposed method is validated across multiple databases, including LivDet 2009, 2011, 2013, and 2015. The quality of the fingerprint samples plays a vital role in PA detection by these methods. Some of the methods including [32], [31] [18] exhibit good performance while being evaluated in the known material scenario but have not been properly tested in the cross-material protocol of FPAD.
2.3 Deep learning-based-methods
Deep Convolutional Neural Networks (CNNs) excel at extracting minute features from input image samples. These models have shown remarkable classification performance on image datasets such as MNIST [7], CIFAR [36], and ImageNet [6]. This ability of the CNNs has prompted researchers to explore their application in the detection of PAs as well. The state-of-the-art methods that use deep CNNs for FPAD are discussed in this section. Park et al. [26] extracted multiple patches from segmented fingerprints and trained a custom-made CNN composed of three convolutional and one fully connected layer. The proposed model utilizes a voting strategy instead of score-level fusion to compute the overall liveness score for an input fingerprint sample. The proposed method is validated using LivDet 2009 database. In a similar attempt, Toosi et al. [34] applied fingerprint foreground segmentation on fingerprint images followed by patch extraction. The extracted patches are further used to train a pre-trained Alexnet [19] CNN classifier. The proposed method is validated using LivDet 2011 and 2013 databases. Nogueira et al. [10] trained two separate SVMs using the features extracted with Local Binary Pattern (LBP) image descriptor and a single-layer CNN respectively. They also incorporated various data augmentation techniques including image rotation and scaling. The limitation of this work is the feature extraction and the classification tasks are performed as separate tasks which do not allow the entire system to learn their parameters together. In another work, Nogueira et al. [24] utilized pre-trained CNN classifiers in their work. They employed VGG, Alexnet, CNN, along with an SVC trained with well-known handcrafted features and applied score-level fusion for the detection of PAs. The proposed method is validated using Livdet 2009, 2011, and 2013 databases. Uliyan et al. [35] proposed a deep feature-based approach for detecting PAs. Their method utilizes a Deep Boltzmann Machine (DBM) for feature extraction and a Restricted Boltzmann Machine (RBM) to establish relationships between these features. Yuan et al. [39] proposed a multi-modal approach utilizing multiple pre-trained CNNs to detect the PAs. The proposed method is validated using LivDet 2011, 2013, 2015, and NUAA face databases. Chugh et al.[5] proposed a method that utilizes minutiae-centered patches for training and testing on the MobileNet classifier. The global liveness score for an input sample is computed by the fusion of the liveness score generated by the classifier for all the patches. The proposed method is validated using LivDet 2011, 2013, 2015, and Michigan State University’s (MSU) FPAD database. Zhang et al. [42] proposed a novel CNN architecture with a series of improved residual connected blocks. It reduces the problem of over-fitting and also takes less processing time. The proposed method is validated on Livdet 2013 and 2015 databases. Since, novel fabrication materials are being discovered nowadays, generalizing an FPAD model to detect PAs in an open-set paradigm is a challenging issue. Chugh et al. [4], in continuation to their previous work [5], proposed an approach for detecting spoofs fabricated with unknown materials. They have presented an image synthesis technique for generating new fingerprint patches which augments the training of MobileNet classifier representing the unknown materials spoofs while experimenting with LivDet 2017, ATVSFFp, and MSU-FPAD databases. Generative Adversarial Networks can also be utilized to augment the training database with spoofs fabricated with unknown materials. In a similar attempt, Rai et al. in [28] utilized Wasserstein GAN to generate fingerprint patches for better training of the DenseNet classifier against the spoofs fabricated with unknown material. The proposed method is validated on LivDet 2015, 2017, and 2019 databases. Although the Deep learning-based methods are quite effective for image classification problems still, they suffer from low classification accuracy in the area of FPAD while being tested in the cross-material protocol of FPAD. The detailed literature survey concludes that the deep learning-based methods are proven to be superior as compared with the methods belonging to other categories but their performance is required to be improved in cross-material protocol of FPAD. This objective can be attained by developing a hybrid model that utilizes the power of both convolutional and handcrafted features. The proposed model should also be efficient enough to classify the input sample in a minimum amount of time to be incorporated with the AFRS.
3 Proposed Work
In this paper, an ensemble of deep and handcrafted features is proposed for the detection of spoofs in intra-sensor same-material and cross-material protocols of FPAD. It consists of two modules that collaborate to detect the PAs. The first module is a DNN that is fed with the features extracted by Local Phase Quantization (LPQ) and existing handcrafted features. The second module is the DenseNet CNN classifier which has shown remarkable performance while being validated on various image databases including imagenet [6]. Both of the modules are fused dynamically which results in better learning of them as compared with their individual performances. The details of LPQ, handcrafted features, DenseNet, and DyFFPAD are given in the following subsections.
3.1 Local Phase Quantization (LPQ)
LPQ [11] works on the Fourier transform’s blur-insensitive property. This makes LPQ robust against the redundant information and blur present in the input sample. We have utilized it as a prominent feature due to its capability of exploiting the minute information that gets missing in the fabrication process of a spoof. The formulation of the LPQ descriptor is denoted as Eq. 1.
$f_{x}(u)=\sum f(y)w(y-x)e^{-j2\pi uy}$ (1)
Here, $f()$ corresponds to the output of the short-term Fourier transform, $f_{x}$ represents the local Fourier coefficients at four distinct frequency values, and $w()$ defines the neighborhood as a window function.
3.2 Handcrafted Features
The fabrication process impacts the quality of the spoofs fabricated with various materials. According to recent research [32] [31], handcrafted features are proven to be useful while detecting PAs. In this work, we have utilized some existing handcrafted features [32], [44] including ridge-valley clarity, gabor quality, number of abnormal ridges and valleys, ridge-valley smoothness, frequency domain analysis, and orientation certainty level to estimate the quality of the input fingerprint sample. The details of these features is provided in the following subsections.
3.2.1 Ridge and Valley Smoothness (RVS and VWS)
It denotes the smoothness of ridge width and valley width which is exhibited by the live sample but not by the spoof. This irregularity in the spoofs is caused because of the varying elasticity levels of spoof fabrication materials. These materials do not possess the elasticity same as human skin. In addition to this, the pressure applied on the fingertip at the time of sample collection is also a potential reason behind this. RWS and VWS are computed block-wise by cropping the rotated block with a vertical ridge-valley structure followed by the binarization. Thereafter pixels of the block are labeled as ridge or valley with the help of a linear regression algorithm. After this, ridge and valley width is computed for each horizontal line of the block. Averaging the standard deviation of ridge and valley width is done to obtain RWS and VWS feature.
3.2.2 Frequency Domain Analysis (FDA)
FDA of a local block is extracted by computing the ridge-valley structure’s 1D signature. The frequency of the sinusoidal ridge-valley structure is calculated using the discrete Fourier transform of this 1D signature. Spoof fingerprint samples exhibit a varying frequency of sinusoidal ridge-valley patterns while it is consistent in live fingerprints.
3.2.3 Orientation Certainty Level (OCL)
OCL is the block-wise intensity of the energy concentration along the prominent ridge direction. The gradient vector is used to calculate the ratio of two eigenvalues of the covariance matrix.
3.2.4 Gabor Quality (G)
A filter bank with several orientations is applied to each pixel to compute gabor quality at the local level. For one or a few filters with orientations comparable to the block orientation, the Gabor response for a fingerprint block with a normal ridge-valley pattern remains high, but it is low and steady for the same block with an improper ridge-valley structure. Finally, the Gabor quality (G) of a block is calculated by taking the standard deviation of the output of the gabor filter bank.
3.2.5 Ridge-Valley Clarity (RVC)
The ridge and valleys present in a block of live fingerprints are found to be separated consistently. This happens due to the uniform elasticity property of human finger skin.This separation, on the other hand, differs in spoof samples due to inappropriate widths of ridges and valleys induced by the different elasticity levels of spoofing materials employed in spoof fabrication. The RVC feature is computed by dividing the original sample into fixed-sized blocks and then computing the average ridge and valley widths in the block.
3.2.6 Number of abnormal ridges and valleys ($R_{ab}$ and $V_{ab}$)
The width of ridges and valleys is found to be in a range of 5 to 10 pixels in samples of human fingers captured from a sensing device with a resolution of 500 dpi. Since the human skin and spoofing materials have different levels of elasticity, the width of ridges and valleys differs in corresponding samples. A ridge or valley in a local block is regarded as abnormal if its width divergence in various rows of the block exceeds a particular threshold.
3.2.7 Feature vector from quality features
The final feature vector which is represented as Eq. (2) is generated using the mean ($\mu$) and standard deviation ($\sigma$) of the features discussed above.
$Q=\{RWS^{\mu},RWS^{\sigma},VWS^{\mu},VWS^{\sigma},R_{ab}^{\mu},V_{ab}^{\mu},RVC^{\mu},$ (2)
$RVC^{\sigma},FDA^{\mu},FDA^{\sigma},OCL^{\mu},OCL^{\sigma},Gabor^{\mu}\}$
3.3 DenseNet-121
DenseNet[15] consists of convolutional layers connected to their previous layers in a feed-forward manner. The majority of CNN designs suffer from the vanishing gradient problem, which occurs when the Rectified Linear Unit (ReLU) is used as an activation function. In deeper CNNs, the gradient of the loss function approaches zero due to the use of ReLU activation functions. This phenomenon further impacts the learning process which results in overfitting of the model. As fingerprint images have limited colour and texture information, typical CNNs encounter this issue when they are utilized in this domain. This problem of vanishing gradient is solved by DenseNet which has skip connections among the layers inside convolutional blocks as shown in Fig. 2.
Figure 2: Internal architecture of DenseNet-121 classifier
(a) Dense Block                                                  (b) Transition Block
Figure 3: Composition of (a) dense block and (b) transition block
DenseNet-121 comprises four dense blocks containing six, twelve, twenty-four, and sixteen convolutional layers, respectively. The internal architecture of DenseNet is illustrated in Figure 2. Each dense block performs three operations: batch normalization, ReLU activation and convolution, and aggregates feature maps from all its preceding layers. The output of a dense block is formulated as shown in Eq. (3).
$X_{n}=F_{n}[A_{0},A_{1},A_{2},A_{3},.....A_{n_{-}1}]$ (3)
where $A_{0},A_{1},A_{n-1}$ represents the concatenation of all feature maps from the previous layers and $F_{n}$ denotes a function that performs batch normalization followed by a convolution operation. In addition, after each dense block, a transition block is utilized to execute a convolution operation with a kernel size of $1\times 1$. Following this convolution procedure, a pooling operator minimizes the size of feature maps after each dense block. Figure 3 depicts the internal architecture of the dense layer and transition layer.
3.4 Deep Neural Network (DNN)
DNN is a type of artificial neural network that has several fully connected layers between the input and output layers. The working of DNNs is inspired by the human brain as its neurons resemble similar characteristics as neurons of the brain. We have used DNN to test the individual performance of the handcrafted features and LPQ image descriptor and to design the proposed architecture.
3.5 Proposed DyFFPAD Model
The architecture of the proposed DyFFPAD is depicted in Fig. 4. It is a dynamic ensemble of handcrafted and convolutional features. The DNN works on the feature vectors generated from handcrafted features and LPQ image descriptors while the convolution base of DenseNet extracts convolutional features from the input samples. The Features extracted from both modules are combined to form a feature vector which is further fed as an input to a DNN that consists of fully connected layers with 48, 16, and 1 neuron respectively. In the forward pass of the training process, a confidence score is produced by the model. The loss of the output score with respect to the expected output is calculated which is back-propagated to the entire model including CNN as well as DNNs for the learning of their parameters. The proposed model works in an end-to-end manner for the forward and backward pass of the training process and reduces the complexities of the ensemble of handcrafted features and convolutional features for the accomplishment of a classification task.
Figure 4: Block diagram of DyFFPAD architecture
3.5.1 Pre-processing of input samples
The fingerprint samples captured with the sensing devices have white space around the fingertip impression. This white space is meaningless and is required to be removed for the extraction of the desired features from fingerprint samples. The input fingerprint undergoes the pre-processing operation for the extraction of the region of interest from the input fingerprint sample.
3.5.2 Feature Extraction
The LPQ and handcrafted features are extracted from the pre-processed fingerprints to form a combined feature vector. The details of the utilized features are described in sections 3.1 and 3.2.
3.5.3 Training of DyFFPAD
The proposed DyFFPAD model is trained from scratch on benchmark LivDet databases. To expedite the training process, we initialized the parameters of DenseNet’s convolution base with ImageNet weights instead of random values. The results of the proposed model under intra-sensor, same-material, and cross-material protocols are discussed in Section 4.
4 Experimental Setup
4.1 Database
Experiments are conducted on the LivDet 2015, 2017, and 2019 databases to validate the performance of the proposed model. Each database includes multiple sensing devices, with fingerprint samples from these devices organized into separate datasets for training and testing. The details of utilized databases including the name of sensing devices, number of training and testing samples, and names of the fabrication materials are mentioned in Table I.
TABLE I: Details of the databases utilized for the validation of DyFFPAD
Database Sensing device Spoof (Train/Test) Live (Train/Test) Spoofing Materials
LivDet 2015 [23] Crossmatch 1473/1448 1000/1500 Ecoflex, Body Double, Playdoh, OOMOO, Gelatine
Greenbit 1000/1500 1000/1000
Digital Persona 1000/1500 1000/1000 Gelatine, Woodglue, Ecoflex, Latex, Liquid Ecoflex, RTV
Biometrika 1000/1500 1000/1000
LivDet 2017 [38] Greenbit 1200/2040 1000/1700 Gelatine, Latex, Liquid Ecoflex, Body Double, Ecoflex, Woodglue
Orcanthus 1180/2018 1000/1700
Digital Persona 1199/2028 999/1692
LivDet 2019 [25] Greenbit 1200/1224 1000/1020 Woodglue, Mix1, Mix2, Liquid Ecoflex , Body Double, Ecoflex
Orcanthus 1200/1088 1000/990 Mix1, Mix2, Liquid Ecoflex, Body Double, Ecoflex, Woodglue
Digital Persona 1000/1224 1000/1099 Mix1, Mix2, Liquid Ecoflex, Ecoflex, Gelatine, Woodglue, Latex
4.2 Performance Metrics
The performance of the proposed model is assessed using the metrics specified by ISO/IEC IS 30107 [1]. The Attack Presentation Classification Error Rate (APCER) measures the proportion of spoof fingerprint samples that are misclassified. The proportion of real fingerprint samples that are misclassified as spoofs is measured by the Bonafide Presentation Classification Error Rate (BPCER). Eq. (4) and Eq. (5) denote APCER and BPCER, respectively.
$APCER=\frac{\textnormal{\# Mis-classified spoof samples}}{\textnormal{Total fake samples}}\times{100}$ (4)
$BPCER=\frac{\textnormal{\# Mis-classified live samples}}{\textnormal{Total live samples}}\times{100}$ (5)
The Average Classification Error (ACE) is determined by averaging APCER and BPCER and is used to assess the overall performance of the FPAD system. The formulation of ACE is represented by Eq. (6).
$ACE=\frac{APCER+BPCER}{2}$ (6)
The ACE is subsequently used to determine the proposed model’s accuracy which is formulated as Eq. (7).
$Accuracy=100-ACE$ (7)
We have also reported the Equal Error Rate (EER) to measure the integrity of our proposed FPAD model. EER is observed by finding an appropriate threshold value where both the APCER and BPCER values are equal.
4.3 Implementation Details
The proposed model is implemented in Python and utilizes the TensorFlow-Keras library. The LPQ and handcrafted features are extracted using Matlab. The NVIDIA TESLA P100 GPU was used for all training and testing. Each model is trained for 300 epochs which requires 4 to 5 hours for its training.
4.4 Ablation Study
Figure 5: Receiver Operating Characteristic (ROC) curves for LivDet 2017 Orcanthus, Digital Persona, and Greenbit sensors
In order to compare the performance of the proposed model with handcrafted features, LPQ image descriptor, and DenseNet classifier, we have trained the corresponding DNN and CNN models. The handcrafted features and LPQ obtained from the fingerprint samples are used for the training of DNN while DenseNet-121 is trained on original fingerprint samples. The performance of the individual models and the proposed model is validated using LivDet 2017 which is a challenging database since it has training and testing sets containing spoof samples from different materials. The findings of this experiment are reported in Table II. Table II clearly indicates the supremacy of DyFFPAD over the DNN trained with LPQ, DNN trained with handcrafted features and LPQ, and the DenseNet classifier. The dynamic ensemble of handcrafted and deep features empowers the model to attain better classification accuracy irrespective of the materials used for the fabrication of spoofs. The proposed DyFFPAD models attain an overall classification accuracy of 96.55% as compared with 79.12% (LPQ+DNN), 75.93% (Handcrafted+DNN), 90.91% (LPQ+Handcrafted+DNN) and 93.64% (DenseNet-121 classifier). The performance comparison of the aforementioned models with DyFFPAD is also shown using the Receiver Operating Characteristics (ROC) curve which is depicted in Fig. 5. ROC curve is a representation of variation in True Positive Rate (TPR) and False Positive Rate (FPR) and is essential in measuring the performance of a machine/deep learning model. The Area Under the Curve (AUC) represents the region enclosed by the plot of the FPR versus TPR. A larger AUC indicates the better proficiency of the model while classifying the input samples.
5 Experimental Results and Comparative Analysis
5.1 Experimental Results
The performance of the proposed method has been validated in two benchmark scenarios: intra-sensor and known spoofing material, and intra-sensor and unknown spoofing material. The following subsections provide detailed descriptions of these scenarios and the corresponding results of the proposed method.
TABLE II: Findings of the ablation study performed on LivDet 2017 database
Method
Accuracy
Accuracy
(Digital Persona)
Accuracy
Average
LPQ+DNN 84.91 77.06 75.40 79.12
Handcrafted+DNN 75 76.62 76.19 75.93
Handcrafted+LPQ+DNN 88.89 90.23 93.62 90.91
DenseNet-121 Classifier 92.79 94.19 93.95 93.64
DyFFPAD 96.34 97.14 96.00 96.49
TABLE III: Performance on LivDet 2015 database
Database Sensor BPCER
APCER
APCER
ACE (%) EER (%)
LivDet 2015 Digital Persona 8.1 4.25 5.74 6.08 6.2
Crossmatch 0.83 0.50 0.68 0.72 0.8
Biometrika 4.5 2.5 3.62 3.52 3.6
Greenbit 9.75 1.9 3.1 5.3 7.9
Average 5.79 2.28 3.28 3.90 4.57
5.1.1 Intra-Sensor and Known Spoofing Material
The training and testing fingerprint samples are acquired using the same sensing device in this experimental setup. Moreover, the spoof samples in training and testing datasets are fabricated using the same materials. LivDet 2015 partially follows this protocol as it consists of two-thirds of the testing spoofs fabricated with known materials and one-third with unknown materials. The results of the proposed model on the LivDet 2015 database are presented in Table III. The table indicates that the proposed model achieves an average BPCER of 5.79% and an APCER of 2.28%, as shown in the column labeled ”APCER (Known).”
TABLE IV: Performance on LivDet 2017 and 2019 databases
Database Sensor BPCER APCER ACE (%) EER (%)
LivDet 2017 Orcanthus 5.59 2.04 3.66 4.2
Digital Persona 3.34 2.46 2.87 2.7
Greenbit 4.68 3.44 4.00 3.9
Average 4.53 2.64 3.51 3.6
LivDet 2019 Digital Persona 12.86 4.74 8.43 9.5
Greenbit 2.06 1.80 1.92 1.8
Orcanthus 2.83 0.65 1.69 1.8
Average 5.91 2.39 4.01 4.23
5.1.2 Intra-Sensor and Unknown Spoofing Material
In this experimental protocol, the training and testing fingerprint samples are acquired using the same sensing device. However, the materials used for the fabrication of training and testing spoof samples are different. This validation protocol is designed to evaluate the FPAD system’s efficacy in a real-attack scenario, wherein an adversary may attempt to present a forged fingerprint artifact created with novel fabrication materials that were not part of the FPAD model’s training data. LivDet 2017 and 2019 are prepared as per this protocol of FPAD. The results of our proposed method, as evaluated on the databases mentioned, are summarized in Table IV. As depicted in Table IV, our proposed model achieves a BPCER of 4.53%, APCER of 2.64%, and an ACE of 3.5%, and EER of 3.6% on the LivDet 2017 database. Similarly, on the LivDet 2019 database, the proposed model showcases the ability to discriminate between live and spoof samples, with a BPCER of 5.91%, APCER of 2.39%, and EER of 4.23%. Notably, when confronted with the spoof samples contained in the LivDet 2015 database, the proposed method exhibits an average APCER of 3.28%, as indicated in the column labeled ”APCER (unknown)” in Table III. It also shows that the proposed methods achieve an ACE of 3.90% and EER of 4.57% on LivDet 2015 database. The live and spoof fingerprint samples exhibit different textures and ridge valley widths due to significant differences between the properties of finger skin and spoofing materials. The convolutional layers of CNNs allow for effective classification by extracting distinguishing features from spoof samples, regardless of the materials used for their fabrication. Moreover, the handcrafted features and image descriptor extract discriminating features for the training of DNN. The dynamic ensemble of DNN and CNN facilitates improved parameter learning for both networks. The proposed architecture reduces the need for training two different models. The findings of the proposed method are compared to existing methods validated on benchmark databases, as discussed in section 5.2.
5.2 Comparative Analysis
In this section, findings of the proposed method are compared with state-of-the-art methods as per the spoofs fabricated with known and unknown materials. A detailed comparative analysis is provided in the following subsections.
TABLE V: Comparison with state-of-the-art methods on LivDet 2015
Method
Accuracy
Accuracy
Accuracy
(Digital Persona)
Accuracy
Average
Kim et al. [18] - - - - 86.39
Uliyan et al. [35] 95.00 - - - 95.00
Sharma et al. [32] 98.07 95.7 94.16 95.22 95.78
Jung et al. [16] 98.60 96.20 90.50 95.80 95.27
Sharma et al. [31] 95.0 96.7 97.2 95.50 96.10
Xia et al.[37] 89.18 95.47 86.28 90.36 90.32
Spinoulas et al. [33] 98.10 98.56 94.80 96.80 97.11
LivDet 2015 Winner [23] 98.10 95.40 93.72 94.36 95.39
Yuan et al.[41] 96.54 95.32 93.2 93.76 94.70
Zhang et al. [43] 97.01 97.81 95.42 97.02 96.82
DyFFPAD 99.29 94.71 93.92 96.48 96.10
5.2.1 Comparison with state-of-the-art methods on LivDet 2015 database
The LivDet 2015 database includes spoof samples fabricated using both known and unknown spoofing materials. Table V provides a detailed comparison of the proposed method’s performance with state-of-the-art methods validated on this database. Table V clearly shows that the proposed method outperforms the methods discussed in [32], [17], [23], [35], [31], [37], [39], and [18]. while it is comparable to the methods discussed in [33] and [43]. The proposed method achieves an accuracy of 99.29% which is the best as compared with all other methods validated on crossmatch sensor. The proposed method outperforms most other methods, achieving an overall classification accuracy of 96.10
5.2.2 Comparison with state-of-the-art methods on LivDet 2017 database
The performance of the proposed method is also compared with methods validated on the LivDet 2017 database. Notably, the training and testing spoof samples within this database are fabricated using different spoofing materials. Our method successfully demonstrates robust classification capabilities in distinguishing between live and spoof samples, achieving commendable accuracy. Table VI shows that the proposed method outperforms the methods discussed in [5], [4], [43], and [12] when evaluated on the fingerprint samples captured with orcanthus and digital persona sensors while it is comparable with [28] on orcanthus sensor. The proposed method achieves an accuracy of 96.00% on greenbit sensor which is better than of [12], [43] and is comparable with [4]. The proposed method also outperforms the state-of-the-art methods in terms of overall classification accuracy of 96.49%. This comparison reveals that the dynamic ensemble of handcrafted and convolutional features is able to detect spoofs regardless of the material used for fabrication.
TABLE VI: Comparison with state-of-the-art methods on LivDet 2017 database
Method
Accuracy
Accuracy
(Digital Persona)
Accuracy
Average
Zhang et al. [43] 93.93 92.89 95.20 94.00
Rai et al. [28] 96.53 93.64 94.74 94.97
Chugh et al. [5] 95.01 95.20 97.42 95.88
Gonzalez et al. [12] 94.38 95.08 94.54 94.66
Chugh et al. [4] 94.51 95.12 96.68 95.43
DyFFPAD 96.34 97.14 96.00 96.49
5.2.3 Comparison with state-of-the-art methods on LivDet 2019 database
Table VII compares the performance of the proposed model with state-of-the-art methods evaluated on the LivDet 2019 database. This comparison reveals that the proposed method achieves an accuracy of 91.57%, outperforming the approaches outlined in [4] as well as the FPAD algorithms, including JungCNN, JWL LivDet, and ZJUT DET, when tested on samples captured with digital persona sensor. For the orcanthus sensor, the proposed method outperforms the majority of the methods with a classification accuracy of 98.32%. The overall performance comparison indicates that the proposed method outperforms all other methods, achieving a classification accuracy of 95.99
The comparative analysis mentioned above concludes that the proposed method consistently outperforms state-of-the-art methods in the intra-sensor paradigm of FPAD, regardless of whether the spoof samples are fabricated with known or unknown materials. As compared to standard CNN-based techniques, the dynamic ensemble of DNN empowered with handcrafted features and CNN enables them to learn their parameters better.
Figure 6: Detection Error Trade-off (DET) curves for LivDet 2015, 2017, and 2019 databases
TABLE VII: Comparison with state-of-the-art methods on LivDet 2019 database
Method
Accuracy
Accuracy
(Digital Persona)
Accuracy
Average
Chugh et al. [4] 97.50 83.64 99.73 93.62
Jung CNN [25] 99.13 81.23 99.06 93.14
JWL LivDet [25] 97.45 88.86 99.20 95.17
ZJUT Det A [25] 97.50 88.77 99.20 95.16
DyFFPAD 98.32 91.57 98.08 95.99
5.3 Evaluation of DyFFPAD in High-Security Systems
In the context of high-security systems, it is imperative to evaluate the performance of an FPAD model comprehensively. While the model’s fundamental objective is indeed to minimize the APCER, BPCER, and ACE, its assessment should extend beyond these metrics to ensure its efficacy in safeguarding high-security environments. In this study, we present the results of the proposed model using the Detection Error Trade-off (DET) curve. The DET curve is a graphical representation of the error rates achieved by a binary classification system as the classification threshold value is varied. The DET curves for all datasets from the LivDet 2015, 2017, and 2019 databases are shown in Fig. 6. As shown in Fig. 6 it is evident that the proposed model achieves a BPCER of less than 1% to achieve APCER of 1% when evaluated on crossmatch, and it is in the range of 15% - 40% for biometrika, greenbit and digital persona sensors of LivDet 2015 database. On the LivDet 2017 database, the model is able to restrict BPCER in the range of 7% - 18% when testing spoof samples are obtained using unknown spoof materials. Similarly, the model maintains a BPCER of less than 5% for greenbit and orcanthus sensors of LivDet 2019 database.
6 Conclusion
In this paper, a novel end-to-end method is presented which utilizes the handcrafted and convolutional features together for the detection of PAs. The increasing deployment of AFRS in security and commercial applications makes them vulnerable to various possible PAs. The proposed method has demonstrated the capability to detect PAs fabricated with known and unknown fabrication materials. Additionally, it suggests a novel ensemble of deep and handcrafted features in an end-to-end manner, making it appropriate for real-time collaboration with the AFRS. The proposed method is tested on fingerprint databases which are prepared as per benchmark experimental protocols. In the future, we will investigate the proposed model’s capabilities for cross-sensor validation using benchmark fingerprint databases.
Declaration
  • The publicly available databases used for the validation of this work can be accessed via the following link: [https://livdet.org/registration.php].
  • Conflict of Interest: The authors declare that there are no conflicts of interest.